OpenStudy (anonymous):

Samantha wants to use her savings of $1150 to buy shirts and watches for her family. The total price of the shirts she bought was $84. The watches cost $99 each. What is the maximum number of watches that Samantha can buy with her savings?

OpenStudy (anonymous):

Let's say the number of watches = x 1150 = 84 + (99x) 1150 - 84 = 99x (1150 - 84)/99 = x

OpenStudy (anonymous):

Round the number of watches down if it's a decimal number. You can't have part of a watch.

OpenStudy (anonymous):

So the answer is 10?

OpenStudy (anonymous):

[[1150\le 99x +84\] solve for x number of watches

OpenStudy (anonymous):

Yeahhh it is.

OpenStudy (anonymous):

Okay, cool. Thanks.